OpenStudy (anonymous):

f(x) = (11 − x^3)^7 find the inverse of this function PLEASE HELP

OpenStudy (anonymous):

switch \(x\) and \(y\) \[x=(11-y^3)^7\] solve for \(y\)

OpenStudy (anonymous):

I did that but I coudnt get why for some reason, my answer is wrong maybe im doing it wrong

OpenStudy (anonymous):

take the seventh root \[\sqrt[7]{x}=11-y^3\] subract 11 \[\sqrt[7]{x}-11=-y^3\] change the sign \[11-\sqrt[7]{x}-11=y^3\] take the cubed root \[\sqrt[3]{11-\sqrt[7]{x}}=y\]

OpenStudy (anonymous):

Thats what I got but my homework website is marking it wrong